OpenStudy (anonymous):

Simplify the radical expression: 3√75

OpenStudy (xapproachesinfinity):

how can we rewrite \(\normalsize\color{green}{ \sqrt{75}\\ }\) like \(\normalsize\color{green}{ \sqrt{8}=\sqrt{4.2}\ = 2\sqrt{2}\ }\)

OpenStudy (xapproachesinfinity):

@SiMPRO

OpenStudy (xapproachesinfinity):

think about the factors of 75?

OpenStudy (xapproachesinfinity):

\(\normalsize\color{maroon}{ \ {1, 3, 5, 15. 25. 75}\ }\)